OpenStudy (anonymous):

predict the number of covalent bonds that the following atoms will form C,H, O, N

OpenStudy (australopithecus):

C = 4 possible bonds H = 1 possible bond O = 2 possible bonds N = 3 to 4 possible bonds

OpenStudy (anonymous):

Thank you, do you mind helping me with another that i need help on

OpenStudy (anonymous):

why are hydrogen and van der waals forces considered weak bonds?

OpenStudy (australopithecus):

because they are constantly being broken and reformed compared to covalent bonds which tend to remain stable

OpenStudy (australopithecus):

it take less energy to break intermolecular bonds than it does to break covalent bond thus they are considered weak

OpenStudy (anonymous):

Such a live saver, I have two more questions that I need clarification on.

OpenStudy (anonymous):

what is a simplified way to explain hydrogen bonding between water molecules?

OpenStudy (australopithecus):

I have no idea what it means by a simplified way to explain hydrogen bonding between water molecules, here is an old note of mine on intermolecular forces

OpenStudy (australopithecus):

hydrogen bonding between water molecules forms a strong network

OpenStudy (australopithecus):

due to water having very unique hydrogen bond angles they can form very strong hydrogen bond networks

OpenStudy (anonymous):

how great a difference is there in hydrogen ion concentration between a ph of 6 and a ph of 8?

OpenStudy (australopithecus):

the lower you go on the pH scale the higher the Hydrogen ion concentration, the higher you go the lower the concentration of hydrogen ions Use the formula pH = - log [ H+]

OpenStudy (australopithecus):

I'm pretty sure the concentration is in molarity (M)

OpenStudy (australopithecus):

do you know how to use the equation? It is pretty easy to solve

OpenStudy (anonymous):

no

OpenStudy (australopithecus):

6 = - log [ H+] -6 = log [ H+] 10^-6 = 10^log[H+] 10^-6 = H+

OpenStudy (australopithecus):

you should be able to calculate the concentration of H+ at 8pH by following my examle

OpenStudy (australopithecus):

Remember 10^log(x) = x

OpenStudy (australopithecus):

\[b^{\log_b(x)} = x\]

OpenStudy (australopithecus):

that is a general rule

OpenStudy (australopithecus):

so at 6pH it has a concentration of 0.000001M

OpenStudy (australopithecus):

do you know how to calculate the molarity now of hydrogen at 8pH?

OpenStudy (australopithecus):

once you do just subtract the two concentrations that will give you the difference.

OpenStudy (anonymous):

a bit lost

OpenStudy (australopithecus):

you dont understand the math?

OpenStudy (australopithecus):

\[\log(x) = \log_{10}(x)\] we just use use log(x) as short hand for log base 10

OpenStudy (australopithecus):

the same rule applies \[10^{\log(x)}\] is the same thing as x

OpenStudy (australopithecus):

if you told me where you were lost I could be of more help but meh